OpenStudy (anonymous):

In an experiment, a ball is drawn from an urn containing 7 orange balls and 12 red balls. If the ball is orange, three coins are tossed. Otherwise two coins are tossed. How many elements of the sample space will have a orange ball ?

OpenStudy (dumbcow):

8 possible "3-coin" combinations ---> 8*7 = 56

OpenStudy (anonymous):

How did you get this?

OpenStudy (dumbcow):

Each time ball is orange (which occurs 7 of 19 possible times) 3 coins are tossed resulting in 8 possible cases HHH HHT HTH HTT THH THT TTH TTT

OpenStudy (anonymous):

okay thanks and How many elements of the sample space are there altogether ?

OpenStudy (anonymous):

Is it 56+8*12 = 56+96 = 152?

OpenStudy (anonymous):

Please reply @dumbcow

OpenStudy (dumbcow):

no because only 2 coins are tossed when ball is red how many combinations do you get when 2 coins are tossed?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So it is 56+12*4 = 56+48 right?

OpenStudy (anonymous):

So 104 @dumbcow right?

OpenStudy (dumbcow):

yep :)